OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos^2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

we have a quadratic equation ....

OpenStudy (amistre64):

how do we solve a quadratic equation?

OpenStudy (amistre64):

if we let u = cos(x) \[cos^2x + 2 ~cos x + 1 = 0\] \[u^2 + 2u + 1 = 0\]

OpenStudy (anonymous):

\[u^2+2u+1=0\] \[(1+u) (1+u)\]

OpenStudy (anonymous):

@amistre64 where do I go from here?

OpenStudy (anonymous):

hmmmmm

OpenStudy (anonymous):

If (u+1)^2 or (u+1)*(u+1) = 0, what solutions do you have for u ? What can you tell about at least one of those two numbers if you know that when they are multiplied with each other they equal 0 ? What can you tell about A (or B) if you know that A * B = 0 ?

OpenStudy (anonymous):

This is kinda hard.....

OpenStudy (anonymous):

WHat grade are you in @dglicks

OpenStudy (anonymous):

@AngusV if i know that A*B=0 then one of those variables is 0

OpenStudy (anonymous):

Perfect. In this case both variables are equal to each other ( so it's kinda like A*A=0 ). Which means A is 0. Now instead of A you have (u+1). Which means (u+1)=0. Which means u=(-1). Recall that u was just a notation (our way of saying) cos(x). That means cos(x)=-1.

OpenStudy (anonymous):

got it thank you @AngusV

OpenStudy (anonymous):

x = 2π x = π x = \[\pi/4, 7 \pi/ 4\] x = \[\pi/2 , 3\pi/2\] @AngusV these are my options

OpenStudy (anonymous):

The only "place" on the trigonometric circle where cos(x)= -1 is on the far (furthest) left - at x=180 degrees or x=pi. At pi/2 and 3pi/2 cos(x) = 0 but you need cos(x) to equal -1. Where did you get those values from ?

OpenStudy (anonymous):

Those values are my options for the multiple choice question

OpenStudy (anonymous):

@AngusV

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so its the orginal question \[\cos^2(x)+2\cos(x) + 1 = 0\]

OpenStudy (campbell_st):

@dglicks is this the question?

OpenStudy (anonymous):

No the question is: Find all solutions in the interval: \[(0,2\pi)\] \[\cos^2x + 2 \cos x + 1 = 0 \]

OpenStudy (anonymous):

So yes that is PART of the question

OpenStudy (campbell_st):

ok... so this is a quadratic equation that can be factored... its a perfect square you used the substitution method and came up with \[(1 + u)(1+u) = (u + 1)^2\] if you reverse the subsitution you get \[(\cos(x) + 1)^2 = 0\] so now you need to solve cos(x) + 1 = 0 for cos(x) can you do that..?

OpenStudy (anonymous):

Im not exactly sure what to do in order to solve for cos(x) + 1= 0

OpenStudy (campbell_st):

well get cos(x) on its own by subtracting 1 from both sides of the equation

OpenStudy (anonymous):

cos(x)= -1 I know how to do that, and to solve for this do I simply put that in my calculator?

OpenStudy (campbell_st):

well remember you are working in radians you can also use the cos curve